OpenStudy (anonymous):

The temperature at 6:00 a.m. was –12°F. By 3:00 p.m. the temperature had risen 21°F. What was the temperature at 3:00 p.m.? A. –33°F B. –9°F C. 9°F D. 33°F Question Resources endlesslight254

OpenStudy (anonymous):

@endlesslight254

OpenStudy (michele_laino):

please what is -12 + 21=...?

OpenStudy (anonymous):

9

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Thx

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

awww u said -12

OpenStudy (anonymous):

i knew it was 9 :D

OpenStudy (anonymous):

12.6 + (–12.6) A. 25.2 B. 24.2 C. 0 D. –25

OpenStudy (anonymous):

A :3 if people r asking for answers why not give it to them

OpenStudy (michele_laino):

please note that I can re-write your computation, as follows: \[12.6 + ( - 12.6) = 12.6 - 12.6 = ...?\] since I applied this rule: \[ + * - = - \]